# Cross-Dataset Gaze Estimation by Evidential Inter-intra Fusion

### v1

@@ Abstract.
Achieving accurate and reliable gaze predictions in complex and diverse environments remains challenging. Fortunately, it is straightforward to access diverse gaze datasets in real-world applications. We discover that training these datasets jointly can significantly improve the generalization of gaze estimation, which is overlooked in previous works. However, due to the inherent distribution shift across different datasets, simply mixing multiple dataset decreases the performance in the original domain despite gaining better generalization abilities. To address the problem of “cross-dataset gaze estimation”, we propose a novel Evidential Inter-intra Fusion (EIF) framework, for training a cross-dataset model that performs well across all source and unseen domains. Specifically, we build independent single-dataset branches for various datasets where the data space is partitioned into overlapping subspaces within each dataset for local regression, and further create a cross-dataset branch to integrate the generalizable features from single-dataset branches. Furthermore, evidential regressors based on the Normal and Inverse-Gamma (NIG) distribution are designed to additionally provide uncertainty estimation apart from predicting gaze. Building upon this foundation, our proposed framework achieves both intra-evidential fusion among multiple local regressors within each dataset and inter-evidential fusion among multiple branches by Mixture of Normal Inverse-Gamma (MoNIG) distribution. Experiments demonstrate that our method consistently achieves notable improvements in both source domains and unseen domains.
Gaze Estimation, Cross-Dataset Training, Evidential Fusion
ccs: Human-centered computing User models
1. Introduction
@@
Gaze estimation is an important task in computer vision used to determine where a person is looking based on visual cues. It has gained attention for its wide range of applications in human-computer interaction (Majaranta and Bulling, 2014; Rahal and Fiedler, 2019), virtual reality (Patney et al., 2016; Kim et al., 2019), and assistive technology (Jiang and Zhao, 2017; Liu et al., 2016; Dias et al., 2020).
@@
Figure 1. Comparison between traditional Cross-Domain scenario and our Cross-Dataset Training scenario. Here, $\mathcal{D}_{s}$ represents the source domain, $\mathcal{D}_{t}$ represents the target domain, and $\mathcal{D}_{u}$ represents the unseen domain. Cross-Dataset Training involves training on multiple source domains with significant distribution discrepancies, aiming for consistent performance across these source domains and the unseen domains.
@@
The challenge of gaze estimation primarily lies in the complexity and diversity of real-world environments. In fact, existing datasets (Funes Mora et al., 2014; Zhang et al., 2017, 2020; Kellnhofer et al., 2019) are collected in different environments, resulting in notable variations in data distribution among them. To enhance the model’s generalization ability across multiple scenarios, current research (Wang et al., 2022a; Guo et al., 2020; Liu et al., 2021; Bao et al., 2022; Cheng et al., 2022; Cai et al., 2023; Liu et al., 2024; Bao and Lu, 2023) focuses on studying cross-domain scenarios and improving performance through single-source domain adaptation and domain generalization strategies, achieving promising progress. However, in practical applications, we often have easy access to multiple data sources, so it is natural to expect models trained on these datasets to perform well across all these source domains as well as unseen domains. This issue is known as cross-dataset training, and Fig. 1 illustrates the differences between our setting and traditional cross-domain setting. While cross-dataset training has been extensively studied in many fields such as object detection (Yao et al., 2020; Chen et al., 2023; Yao et al., 2020), semantic segmentation (Shi et al., 2021; Wang et al., 2022b), facial expression recognition (Zeng et al., 2018) and trajectory prediction (Gilles et al., 2022), it has received little attention in gaze estimation.
@@
Figure 2. Test accuracy on different datasets with varied combination of training data.
@@
Preliminary experiments depicted in Fig. 2 show that simply combining data (e.g., Gaze360+ETH-Gaze, shorted as G+E) from multiple datasets can significantly enhance the model’s performance on unseen domains (e.g., MPIIFaceGaze, EyeDiap), even outperforming carefully designed state-of-the-art (SOTA) single-source domain generalization methods (Cheng et al., 2022; Bao and Lu, 2023). This improvement is due to the diverse data distributions across source domain datasets. By merging these datasets, we increase the variety of the training data, which largely enhances the model’s generalization ability to unseen domains. Unfortunately, despite increasing the volume of training data, cross-dataset training results in diminished performance on test sets of the original source domains (Gaze360, ETH-Gaze) compared to training exclusively with their own training data. This decrease is also attributed to the significant differences in data distribution among source domains, but need to be mitigated.
@@
To prevent the degradation of source domain performance and further enhance performance on unseen domains, in this paper, we propose a novel solution called Evidential Inter-intra Fusion (EIF) for cross-dataset gaze estimation. Acknowledging the inherent differences between the datasets, our approach involves training independent single-dataset branches for each dataset, thereby mitigating source domain performance degradation. Additionally, building upon the findings from preliminary experiments, which suggest that merging rich source domain information benefits unseen domains, we devise a cross-dataset branch containing only the deep layers, where the generalizable features from each source domain branch are integrated, thereby boosting performance on unseen domains. On the other hand, to combat the performance drop in the source domains, we argue that the non-stationary gazing process inherent in each dataset should be carefully considered, where non-stationary refers to our gaze being primarily influenced by head posture and eye movement with their effects varying across different gaze ranges. Such challenge is tackled by proposing to partition the data space into overlapping subspaces, with an independent local regressor assigned to each subspace. Furthermore, in order to select appropriate regressors both in within-dataset and cross-dataset branches, we design to employ evidential regressors based on the Normal and Inverse-Gamma (NIG) distribution for not only predicting gaze but also offering uncertainty estimation. Leveraging the Mixture of Normal Inverse-Gamma (MoNIG) distribution, our framework achieves both intra-evidential fusion among numerous local regressors within each dataset and inter-evidential fusion across multiple-dataset branches.
@@
During training, to enable the model to rapidly adapt to variations in the number of datasets, the training process of our proposed EIF is divided into two stages. In stage 1, we train each single-dataset branch separately. In stage 2, we create a cross-dataset branch and jointly train all branches for only a few epochs. During inference, the predictions including both gaze and uncertainty are output, which largely enhance the reliability of our proposed framework. We thoroughly evaluate the effectiveness of our proposed EIF on established gaze estimation benchmarks.
@@
  • We introduce the task of cross-dataset training for the first time in the field of gaze estimation, which can significantly improve the generalization of gaze estimation.
@@
  • We propose a novel framework, called Evidential Inter-intra Fusion (EIF), which consists of several single-dataset branches and a cross-dataset branch, to tackle the challenge in cross-dataset gaze estimation. EIF is equipped with multiple local regressors fused by intra-evidential within each dataset and meanwhile performs inter-evidential fusion across multiple datasets, ensuring accurate gaze estimation in complex and diverse environments.
@@
  • We conduct comprehensive experiments that demonstrate the consistent performance improvement in both source domains and unseen domains.
2. Related Work
2.1. Gaze Estimation
@@
Gaze estimation methods can be divided into two main types (Hansen and Ji, 2009): model-based and appearance-based approaches. Model-based methods  (Alberto Funes Mora and Odobez, 2014; Valenti et al., 2011) rely on the accurate reconstruction of 3D model of the eyes, thus requiring specialized equipments such as depth sensors, infrared cameras, and lighting. By contrast, appearance-based methods (Cheng et al., 2021) estimate gaze using images captured by a single webcam, and thus facilitate the popularity of gaze estimation. With advancements in deep learning, appearance-based methods have significantly improved the performance and garnered widespread attention. However, deep learning methods heavily rely on data that can ideally reflect the distribution of real-world conditions. Yet, the actual gaze estimation scenario is diverse and complex (Ghosh et al., 2023).
@@
In fact, the current gaze datasets are collected in various environments, resulting in significant differences in data distribution. For example, MPIIFaceGaze (Zhang et al., 2017) is collected during the natural use of the laptops, with images captured using the laptop camera. EyeDiap (Funes Mora et al., 2014), on the other hand, is collected in laboratory settings, where participants are instructed to observe given targets, and images are captured using depth cameras. Gaze360 (Kellnhofer et al., 2019) is collected in the wild using a 360-degree camera to simultaneously record multiple participants. Lastly, ETH-Gaze (Zhang et al., 2020) is constructed in a laboratory green screen environment, where lighting conditions are simulated using controlled lighting, and images are captured using ultra-high-resolution cameras.
@@
Due to variations in data distributions among datasets, traditional deep learning methods often struggle in cross-domain scenarios. Many researches focus on employing domain generalization and domain adaptation strategies to alleviate the effect of domain discrepancy. Liu et al. (Liu et al., 2021) propose outlier-guided collaborative adaptation. Wang et al. (Wang et al., 2022a) introduce regression contrastive learning techniques. Bao et al. (Bao et al., 2022) propose a rotation consistency strategy. Cheng et al. (Cheng et al., 2022) utilize adversarial reconstruction techniques to purify gaze features. Cai et al. (Cai et al., 2023) attempt to reducing sample and model uncertainty. Bao et al. (Bao and Lu, 2023) propose constructing Physics-Consistent Features (PCF) to incorporate physical definition.
@@
Unlike the methods mentioned above, our work introduces a novel practical scenario called cross-dataset training. In this scenario, we have access to data from multiple source domains collected in different environments with distribution variations during training, which significantly improves cross-domain performance without sacrificing performance of within-domain.
2.2. Cross-dataset Training
@@
In many computer vision tasks, researches on cross-dataset training have been well-explored, but they face different challenges compared to “cross-dataset gaze estimation”. For instance, in object detection, the main challenge lies in the disparate definitions and granularity of classes across different datasets. Therefore, the focus of related works (Yao et al., 2020; Chen et al., 2023; Shi et al., 2021) is on integrating diverse datasets into a unified taxonomy. In the field of facial expression recognition, a primary challenge in cross-dataset training stems from the inconsisten annotations among datasets. To address this issue, the method (Zeng et al., 2018) learns a transition matrix for modeling the relationship between the latent ground truth and the dataset annotation.
@@
Unlike the tasks mentioned above, the challenge need to be solved in cross-dataset gaze estimation is the inherent distribution shift across different datasets. Similar challenges have also been identified in semantic segmentation for autonomous driving by Wang et al. (Wang et al., 2022b). They propose a solution based on the characteristics of their task, which involves training independent batch normalization (BN) layers for each dataset while sharing their convolutional layers. However, it requires prior knowledge of which source domain the test sample belongs to for selecting the corresponding BN layer during testing. In contrast, our objective is to enable testing in both the source domains and unseen domains for gaze estimation task without requiring to know which corresponding source domain dataset a test sample belongs to.
t Refer to caption
@@
Figure 3. Illustration of the proposed Evidential Inter-intra Fusion (EIF) framework. The model consists of multiple single-dataset branches and a cross-dataset branch, with each branch comprising multiple local regressors. We utilize the intra-evidential fusion module to perform local regressor fusion within each branch and the inter-evidential fusion module to perform fusion across multiple branches. The output includes both predictions and uncertainty. Due to gaze estimation involving yaw and pitch components, there are two identical regression layers with the same model structure.
3. Methods
@@
In this section, we elaborate the proposed Evidential Inter-intra Fusion (EIF) framework. As depicted in Fig. 3, EIF model incorporates multiple single-dataset branches, each dedicated to a distinct source domain dataset. Additionally, to enhance performance on unseen domains, we introduce a cross-dataset branch that aggregates features from all source domain branches. To address the performance degradation of source domain caused by non-stationary nature of the gazing process, we draw inspiration from similar challenges addressed in other tasks such as age estimation (Li et al., 2019), action quality assessment (Yu et al., 2021), equipping each branch with multiple local regressors tailored to specific data subspaces. Subsequently, we employ intra-evidential fusion module to fuse local regressors within each branch and inter-evidential fusion module to fuse multiple branches. In the following subsections, we will delve into the components of the framework, encompassing the single-dataset branch, cross-dataset branch, evidential fusion module, and the two-stage training process.
3.1. Single-dataset Branch
@@
Due to the inherent distributional shifts among datasets, each single-dataset branch is tailored to a specific source domain dataset. It includes several layers for feature extraction and two regression layers, $R_{1}$ and $R_{2}$, designed to estimate the two components of gaze: yaw ($Y_{1}$) and pitch ($Y_{2}$). Since these layers have the same structure, differing only in the gaze components they estimate, we will use $R$ for the regression layers and $Y$ for the gaze components to introduce the model structure conveniently in the following description.
@@
Recognizing the non-stationary nature of gaze behavior might lead to the performance drop when combining multiple datasets, we design the regression layer $R$ to include multiple local regressors, denoted as $R=\{r_{1},r_{2},\ldots,r_{G}\}$. These local regressors operate exclusively within specific data subspaces, which are partitioned based on gaze labels $Y$, and thus can be viewed as a collective of experts, with each expert specializing in specific regression scenarios.
@@
Moving forward, the method of obtaining data subspaces by gaze label $R$ is crucial. Since gaze labels are not uniformly distributed, simply dividing the entire range equitably may result in an imbalanced data distribution among the subspaces. To tackle this issue, we introduce a density-based partitioning approach. Initially, we sort the gaze labels $Y=\{y^{1},y^{2},\ldots,y^{S}\}$ in ascending order to achieve $Y^{*}=\{y^{1^{*}},y^{2^{*}},\ldots,y^{S^{*}}\}$. Subsequently, we sample indices at regular intervals to identify the centers of each gaze group:
@@
(1) $\zeta_{\text{center}}^{g}=Y^{*}(\left\lfloor S\times\frac{(g+0.5)}{G}\right\rfloor)$
[…]
@@
Additionally, we employ a dense overlap strategy among gaze groups, enabling adjacent groups to have overlapping intervals. This strategy ensures that data within overlapping regions contribute to training multiple regressors, thereby enhancing model robustness and accuracy across varying gaze behaviors. The degree of overlap between them is determined by a dense overlap coefficient denoted by $\alpha$. Formally, the boundaries for each group are defined as follows:
@@
(2) $\begin{split}\zeta_{\text{left}}^{g}&=Y^{*}\left(\left\lfloor S\times\frac{(g+0.5-0.5\times\alpha)}{G}\right\rfloor\right),\\ \zeta_{\text{right}}^{g}&=Y^{*}\left(\left\lfloor S\times\frac{(g+0.5+0.5\times\alpha)}{G}\right\rfloor\right),\\ \mathcal{I}^{g}&=(\zeta_{\text{left}}^{g},\zeta_{\text{right}}^{g}).\end{split}$
@@
Here, $\mathcal{I}^{g}$ is boundaries of the $g$-th gaze group. We set $\alpha>1$ to ensure that each gaze group has an overlap interval. By partitioning the gaze groups in this manner, data with gaze label $y_{s}$ that fall within $(\zeta_{\text{left}}^{g},\zeta_{\text{right}}^{g})$ are utilized for training the branch $r_{g}$, and the output range of $r_{g}$ is also constrained to $(\zeta_{\text{left}}^{g},\zeta_{\text{right}}^{g})$.
3.2. Cross-dataset Branch

### v2

@@ Abstract
Achieving accurate and reliable gaze predictions in complex and diverse environments remains challenging. Fortunately, it is straightforward to access diverse gaze datasets in real-world applications. We discover that training these datasets jointly can significantly improve the generalization of gaze estimation, which is overlooked in previous works. However, due to the inherent distribution shift across different datasets, simply mixing multiple dataset decreases the performance in the original domain despite gaining better generalization abilities. To address the problem of “cross-dataset gaze estimation”, we propose a novel Evidential Inter-intra Fusion (EIF) framework, for training a cross-dataset model that performs well across all source and target domains. Specifically, we build independent single-dataset branches for various datasets where the data space is partitioned into overlapping subspaces within each dataset for local regression, and further create a cross-dataset branch to integrate the generalizable features from single-dataset branches. Furthermore, evidential regressors based on the Normal and Inverse-Gamma (NIG) distribution are designed to additionally provide uncertainty estimation apart from predicting gaze. Building upon this foundation, our proposed framework achieves both intra-evidential fusion among multiple local regressors within each dataset and inter-evidential fusion among multiple branches by Mixture of Normal Inverse-Gamma (MoNIG) distribution. Experiments demonstrate that our method consistently achieves notable improvements in both source domains and target domains.
1 Introduction
@@
Gaze estimation is an important task in computer vision used to determine where a person is looking based on visual cues. It has gained attention for its wide range of applications in human-computer interaction [27, 29], virtual reality [28, 21], and assistive technology [19, 24, 13].
@@
Figure 1: Comparison between traditional Cross-Domain scenario and our Cross-Dataset Training scenario. Here, $\mathcal{D}_{s}$ represents the source domain, $\mathcal{D}_{t}$ represents the target domain. Cross-Dataset Training involves training on multiple source domains with significant distribution discrepancies, aiming for consistent performance across source domains and target domains.
@@
The challenge of gaze estimation primarily lies in the complexity and diversity of real-world environments. In fact, existing datasets [14, 40, 41, 20] are collected in different environments, resulting in notable variations in data distribution among them. To enhance the model’s generalization ability across multiple scenarios, current researches [35, 17, 25, 5, 11, 6, 23, 4] focus on studying cross-domain scenarios and improving performance through single-source domain adaptation and domain generalization strategies, achieving promising progress. However, in practical applications, we often have easy access to multiple data sources, so it is natural to expect models trained on these datasets to perform well across all these source domains as well as target domains. This issue is known as cross-dataset training. Fig. 1 illustrates the differences between our setting and traditional cross-domain setting. While cross-dataset training has been extensively studied in many fields such as object detection [37, 7, 37], semantic segmentation [31, 34], facial expression recognition [39] and trajectory prediction [16], it has received little attention in gaze estimation.
@@
Figure 2: Test accuracy on different datasets with varied combination of training data.
@@
Preliminary experiments depicted in Fig. 2 show that simply combining data (e.g., Gaze360+ETH-Gaze, shorted as G+E) from multiple datasets can significantly enhance the model’s performance on target domains (e.g., MPIIFaceGaze, EyeDiap), even outperforming carefully designed state-of-the-art (SOTA) single-source domain generalization methods [11, 4]. This improvement is due to the diverse data distributions across source domain datasets. By merging these datasets, we increase the variety of the training data, which largely enhances the model’s generalization ability to target domains. Unfortunately, despite increasing the volume of training data, cross-dataset training results in diminished performance on test sets of the original source domains (Gaze360, ETH-Gaze) compared to training exclusively with their own training data. This decrease is also attributed to the significant differences in data distribution among source domains, but need to be mitigated.
@@
To prevent the degradation of source domain performance and further enhance performance on target domains, in this paper, we propose a novel solution called Evidential Inter-intra Fusion (EIF) for cross-dataset gaze estimation. Acknowledging the inherent differences between the datasets, our approach involves training independent single-dataset branches for each dataset, thereby mitigating source domain performance degradation. Additionally, building upon the findings from preliminary experiments, which suggest that merging rich source domain information benefits target domains, we devise a cross-dataset branch containing only the deep layers, where the generalizable features from each source domain branch are integrated, thereby boosting performance on target domains.
On the other hand, to combat the performance drop in the source domains, we argue that the non-stationary gazing process inherent in each dataset should be carefully considered, where non-stationary refers to our gaze being primarily influenced by head posture and eye movement with their effects varying across different gaze ranges. Such challenge is tackled by proposing to partition the data space into overlapping subspaces, with an independent local regressor assigned to each subspace. Furthermore, in order to select appropriate regressors both in within-dataset and cross-dataset branches, we design to employ evidential regressors based on the Normal and Inverse-Gamma (NIG) distribution for not only predicting gaze but also offering uncertainty estimation. Leveraging the Mixture of Normal Inverse-Gamma (MoNIG) distribution, our framework achieves both intra-evidential fusion among numerous local regressors within each dataset and inter-evidential fusion across multiple-dataset branches.
@@
During training, to enable the model to rapidly adapt to variations in dataset combinations, the training process of our proposed EIF is divided into two stages. In stage 1, we train each single-dataset branch separately. In stage 2, we create a cross-dataset branch and jointly train all branches for only a few epochs. During inference, the predictions including both gaze and uncertainty are output, which largely enhance the reliability of our proposed framework. We thoroughly evaluate the effectiveness of our proposed EIF on established gaze estimation benchmarks.
@@
  • We introduce the task of cross-dataset training for the first time in the field of gaze estimation, which can significantly improve the generalization of gaze estimation while avoiding performance degradation in the source domains.
@@
  • We propose a novel framework, called Evidential Inter-intra Fusion (EIF) to tackle the challenge in cross-dataset gaze estimation. EIF is equipped with multiple local regressors fused by intra-evidential within each dataset and meanwhile performs inter-evidential fusion across multiple datasets, ensuring accurate gaze estimation in complex and diverse environments. Additionally, EIF is designed with a two-stage training process, enabling it to quickly adapt to variations in dataset combinations.
@@
  • We conduct comprehensive experiments that demonstrate the consistent performance improvement in both source domains and target domains.
2 Related Work
2.1 Gaze Estimation
@@
Gaze estimation methods can be divided into two main types [18]: model-based and appearance-based approaches. Model-based methods [1, 33] rely on the accurate reconstruction of 3D model of the eyes, thus requiring specialized equipments such as depth sensors, infrared cameras, and lighting. By contrast, appearance-based methods [10] estimate gaze using images captured by a single webcam, and thus facilitate the popularity of gaze estimation. With advancements in deep learning, appearance-based methods have significantly improved the performance and gained widespread attention. However, deep learning methods heavily rely on data that can ideally reflect the distribution of real-world conditions. Yet, the actual gaze estimation scenario is diverse and complex [15]. Specifically, we provide a visual display of the different dataset collection environments and example samples in the supplementary.
@@
In fact, the current gaze datasets are collected in various environments, resulting in significant differences in data distribution. For example, MPIIFaceGaze [40] is collected during the natural use of the laptops, with images captured using the laptop camera. EyeDiap [14], on the other hand, is collected in laboratory settings, where participants are instructed to observe given targets, and images are captured using depth cameras. Gaze360 [20] is collected in the wild using a 360-degree camera to simultaneously record multiple participants. Lastly, ETH-Gaze [41] is constructed in a laboratory green screen environment, where lighting conditions are simulated using controlled lighting, and images are captured using ultra-high-resolution cameras.
@@
Due to variations in data distributions among datasets, traditional deep learning methods often struggle in cross-domain scenarios. Many researches focus on employing domain generalization and domain adaptation strategies to alleviate the effect of domain discrepancy. Liu et al. [25] propose outlier-guided collaborative adaptation. Wang et al. [35] introduce regression contrastive learning techniques. Bao et al. [5] propose a rotation consistency strategy. Cheng et al. [11] utilize adversarial reconstruction techniques to purify gaze features. Cai et al. [6] attempt to reducing sample and model uncertainty. Bao et al. [4] propose constructing Physics-Consistent Features (PCF) to incorporate physical definition. These methods enhance the performance of the target domain through carefully designed strategies, but they may do so at the cost of sacrificing source domain performance [4].
@@
Unlike the methods mentioned above, our work introduces a novel practical scenario called cross-dataset training. In this scenario, we have access to data from multiple source domains collected in different environments with distribution variations during training, which significantly improves target domain performance without sacrificing performance of source domain.
2.2 Cross-dataset Training
@@
In many computer vision tasks, researches on cross-dataset training have been well-explored, but they face different challenges compared to “cross-dataset gaze estimation”. For instance, in object detection, the main challenge lies in the disparate definitions and granularity of classes across different datasets. Therefore, the focus of related works [37, 7, 31] is on integrating diverse datasets into a unified taxonomy. In the field of facial expression recognition, a primary challenge in cross-dataset training stems from the inconsistent annotations among datasets. To address this issue, Zeng et al. [39] learn a transition matrix for modeling the relationship between the latent ground truth and the dataset annotation.
@@
Unlike the tasks mentioned above, the challenge need to be solved in cross-dataset gaze estimation is the inherent distribution shift across different datasets. Similar challenges have also been identified in semantic segmentation for autonomous driving by Wang et al. [34]. They propose a solution based on the characteristics of their task, which involves training independent batch normalization (BN) layers for each dataset while sharing their convolutional layers. However, it requires prior knowledge of which source domain the test sample belongs to for selecting the corresponding BN layer during testing. In contrast, our objective is to enable testing in both source domains and target domains for gaze estimation task without needing to know which specific source domain dataset.
3 Methods
@@
Figure 3: Illustration of the proposed Evidential Inter-intra Fusion (EIF) framework. The model consists of multiple single-dataset branches and a cross-dataset branch, with each branch comprising multiple local regressors. We utilize the intra-evidential fusion module to perform local regressor fusion within each branch and the inter-evidential fusion module to perform fusion across multiple branches. These evidence fusion modules are formulated without any learnable parameters. The output includes both predictions and uncertainty. Due to gaze estimation involving yaw and pitch components, there are two identical regression layers with the same model structure.
@@
In this section, we elaborate the proposed Evidential Inter-intra Fusion (EIF) framework. As depicted in Fig. 3, EIF model incorporates multiple single-dataset branches (Sec. 3.1), each dedicated to a distinct source domain dataset. Additionally, to enhance performance on target domains, we introduce a cross-dataset branch (Sec. 3.2) that aggregates features from all source domain branches. To address the performance degradation of source domain caused by non-stationary nature of the gazing process, we draw inspiration from similar challenges addressed in other tasks such as age estimation [22], action quality assessment [38], equipping each branch with multiple local regressors tailored to specific data subspaces. Subsequently, we employ intra-evidential fusion module (Sec. 3.3) to fuse local regressors within each branch and inter-evidential fusion module (Sec. 3.3) to fuse multiple branches.
3.1 Single-dataset Branch
@@
Due to inherent distributional shifts among datasets, each single-dataset branch is tailored to a specific source domain dataset. It includes several layers for feature extraction and two regression layers, $R_{1}$ and $R_{2}$, designed to estimate the two components of gaze: yaw ($Y_{1}$) and pitch ($Y_{2}$). Since these layers have the same structure, differing only in the gaze components they estimate, we use $R$ for the regression layers and $Y$ for the gaze components to introduce the model structure conveniently in the following description.
@@
Recognizing the non-stationary nature of gaze behavior might lead to the performance drop when combining multiple datasets, we design the regression layer $R$ to include multiple local regressors, denoted as $R=\{r_{1},r_{2},\ldots,r_{G}\}$. The outputs of these local regressors are centered around specific gaze labels $Y$, thereby forming a collective of experts, each specializing in a designated gaze region and corresponding regression scenario.
@@
Moving forward, the method of partitioning gaze label intervals $Y$ is crucial. Given the non-uniform distribution of gaze labels, a straightforward division may lead to imbalanced data in each gaze subspace. To address this, we introduce a density-based partitioning approach. First, we sort the gaze labels $Y=\{y^{1},y^{2},\ldots,y^{S}\}$ in ascending order to produce $Y^{*}=\{y^{1^{*}},y^{2^{*}},\ldots,y^{S^{*}}\}$. We then sample indices at regular intervals to locate the mean position for each gaze group:
@@
$\zeta_{\text{center}}^{g}=Y^{*}\left(\left\lfloor S\times\frac{(g+0.5)}{G}\right\rfloor\right),$ (1)
where $\zeta_{\text{center}}^{g}$ is the center of the $g$-th gaze group, $Y^{*}(s)$ denotes the $s$-th element in $Y^{*}$, $S$ is the total sample count, $g$ is the gaze group index, and $G$ represents the total number of gaze groups.
@@
Additionally, we employ a dense overlap strategy among gaze groups, enabling adjacent groups to have overlapping intervals. This strategy ensures that data within overlapping regions contribute to training multiple regressors, thereby enhancing model robustness and accuracy across varying gaze behaviors. The degree of overlap between them is determined by a dense overlap coefficient denoted by $\alpha$. Formally, the boundaries and length for each group are defined as follows:
@@
$\begin{split}\zeta_{\text{left}}^{g}&=Y^{*}\left(\left\lfloor S\times\frac{(g+0.5-0.5\times\alpha)}{G}\right\rfloor\right),\\ \zeta_{\text{right}}^{g}&=Y^{*}\left(\left\lfloor S\times\frac{(g+0.5+0.5\times\alpha)}{G}\right\rfloor\right),\\ \mathcal{I}^{g}&=(\zeta_{\text{left}}^{g},\zeta_{\text{right}}^{g}),\quad L^{g}=\zeta_{\text{right}}^{g}-\zeta_{\text{left}}^{g},\end{split}$ (2)
@@
where $\mathcal{I}^{g}$ represents the boundaries of the $g$-th gaze group, and $L^{g}$ denotes its length. We set $\alpha>1$ to ensure that each gaze group has an overlap interval. By partitioning the gaze groups in this manner, data with gaze label $y^{s}$ that fall within $(\zeta_{\text{left}}^{g},\zeta_{\text{right}}^{g})$ are utilized for training the branch $r_{g}$.
In this setup, each regressor $r_{g}$ is responsible for a specific gaze group and outputs an offset $\Delta^{g}$. The final prediction is derived by combining this offset with the center $\zeta_{\text{center}}^{g}$ and the interval length $L^{g}$, which results in:
$\delta_{LR_{g}}=\Delta^{g}\cdot\frac{L^{g}}{2}+\zeta_{\text{center}}^{g}.$ (3)
This offset-based approach allows each regressor to capture variations within its designated interval, thus enhancing the model’s adaptability across diverse gaze behaviors.
3.2 Cross-dataset Branch